Lulabalula:

Raj gave his mother a bouquet of 12 pink and white roses. There were 4 pink roses in the bouquet. Which ratio represents the number of white roses to the total number of roses in the bouquet?

Convert:

12-4=8 8/12 simplified= 2/3

Lulabalula:

wait what

Convert:

Because they give you the number of pink rose and all the rose are either pink of white you would subtract them from the total amount of roses

Convert:

Therefore the 12-4 which would give you 8 so 8/12

Lulabalula:

oh I thought you would divide

Convert:

No, I don't believe that's what their asking

Lulabalula:

oh ok